# Scalable Multi-Robot Motion Planning Using Guidance-Informed Hypergraphs

### v3

@@ Abstract
In this work, we propose a method for multiple mobile robot motion planning that efficiently plans for robot teams up to an order of magnitude larger than existing state-of-the-art methods in congested settings with narrow passages in the environment. We achieve this improvement in scalability by adapting the state-of-the-art Decomposable State Space Hypergraph (DaSH) planning framework to expand the set of problems it can support to include those without a highly structured planning space and those with kinodynamic constraints. We accomplish this by exploiting guidance about a problem’s structure to limit exploration of the planning space and through modifying DaSH’s conflict resolution scheme. This guidance captures when coordination between robots is necessary, allowing us to decompose the intractably large multi-robot search space while limiting risk of inter-robot conflicts by composing relevant robot groups together while planning.
@@ I Introduction
Multi-robot systems are becoming increasingly prevalent in settings like warehouses and factories. These often constrained environments require careful yet computationally fast motion planning for each robot to accomplish its task without collision. Some navigation approaches compute individual decoupled paths for each robot and resolve conflicts online [1]; however, in congested environments, this introduces the risk of deadlock. Alternatively, coupled offline multi-robot motion planning (MRMP) approaches [1] plan for all robots together in the composite space, the joint planning space of all robots. While these methods provide the high level of coordination required to solve difficult problems, due to the large size of the search space, they are computationally intractable for large robot teams. Hybrid methods [2, 3, 4] seek to combine the benefits of decoupled and coupled planning, providing an increased level of coordination when necessary while also maintaining scalability.
Figure 1: A comparison of the DaSH framework and the Guided-DaSH framework. In DaSH, the task space hypergraph and motion hypergraph are exhaustively constructed. The motion hypergraph is then queried to find an optimistic schedule which then undergoes conflict resolution. In Guided-DaSH, a heuristic search over the implicit problem space allows us to build a minimal portion of the task space hypergraph and motion hypergraph. Because of this, we can directly extract the motion solution from the motion hypergraph.
The Decomposable State Space Hypergraph (DaSH) [5] framework presents a general structure for hybrid planning that benefits from a sparse representation of the planning space. This representation models when coordination between robots is needed, allowing the planner to stay in low-dimensional search spaces when possible, and captures the connections between these spaces. The framework features a hierarchical approach consisting of a high-level representation of the planning space, the task space hypergraph, that directs construction of a motion hypergraph from which an optimistic motion solution which is not guaranteed to be collision-free is found and then validated. The DaSH framework can be applied to develop methods for a variety of multi-robot planning problems. Here, we propose leveraging DaSH for problems including mobile robot motion planning in environments with narrow passages, a classical challenge in motion planning exacerbated in the multi-robot setting, where the robots may also have kinodynamic constraints which further complicate planning.
@@
DaSH proposes an approach to exhaustively constructing the task space hypergraph that performs well for problems like multi-manipulator rearrangement [5], where the search space has an exploitable structure since actions can be classified as pick, place, handoff, etc., but becomes intractable for problems which lack this highly structured search space, such as mobile robot motion planning. Our insight is that we can leverage information about the planning problem to induce this structure while building the high-level representation and further exploit this guidance, in conjunction with a modified conflict resolution scheme, to efficiently search the planning space and support kinodyanmic constraints. In the case of mobile robot motion planning, we can effectively use workspace information to provide this guidance [6, 7, 8, 9, 10]. As such, we propose Guided-DaSH, our adaptation to the DaSH framework to support a wider range of problems and Workspace Guided-DaSH, an application of our framework to the problem of multiple mobile robot motion planning in environments with narrow passages. Our contributions include:
@@
  • An adaptation of the DaSH framework to multi-robot planning problems without a highly structured search space and a novel hybrid MRMP method built on our adapted framework that leverages the workspace to guide planning in multiple ways.
@@
  • Support for a broader range of problems, including planning under kinodynamic constraints within our framework via an adapted conflict resolution structure.
@@ II-A Motion Planning Preliminaries
In motion planning, a configuration refers to a set of values for a robot’s degrees of freedom, including their position in the workspace, orientation, and other parameters such as joint angles. The configuration space ($\mathcal{C}_{space}$) consists of all possible configurations and can be partitioned into $\mathcal{C}_{free}$ and $\mathcal{C}_{obst}$ made up of valid and invalid configurations respectively. Motion planning is the problem of finding a path from a start configuration $q_{s}$ to a goal configuration $q_{g}$ through $\mathcal{C}_{free}$. Computing the full $\mathcal{C}_{obst}$ is intractable [11, 12], thus state-of-the-art methods turned to sampling-based motion planning [13, 14]. Two commonly used methods, Rapidly-exploring Random Trees (RRT) [14] and Probabilistic Roadmaps [13] build tree-based and graph-based representations of $\mathcal{C}_{free}$ respectively.
These methods have also been extended to problems which require satisfying kinematic constraints that limit the robot’s position, and dynamic constraints that limit the velocity or acceleration of the robot. Kinodynamic RRT [15] extends RRT to work with kinodynamic constraints through extensions of randomly sampled control inputs for random durations until a goal region is reached.
@@ II-B Multi-robot Motion Planning
The MRMP problem is an extension of the motion planning problem which consists of finding valid paths $\Pi=\{\pi_{r_{1}},...,\pi_{r_{n}}\}$ for a set of robots $R=\{r_{1},...,r_{n}\}$ that also avoid inter-robot collision. Multi-agent pathfinding (MAPF) is the discrete state space equivalent of the MRMP problem which consists of finding collision-free paths for a set of robots over a given graph representation (e.g. a grid) rather than through continuous space. In environments with narrow passages, it is non-trivial to construct a graph representation that captures the connectivity of the planning space, hence our consideration of MRMP.
Some coupled methods, including Composite PRM [1] and Composite RRT [14] provide a high level of coordination by planning directly in the composite space, $C_{0}\times...\times C_{n-1}$ where $C_{i}$ is the configuration space of robot $i$. This becomes computationally intractable due to the high dimensionality of the space with large groups of robots. Other coupled methods, including MRdRRT [16] and its variants [17, 18], construct single-robot roadmaps and then use these to search an implicit composite space roadmap.
Hybrid methods combine decoupled and coupled planning. CBS-MP [2] performs a low-level search to find decoupled paths for each robot and resolves conflicts using a high-level search in the composite space. Several hybrid MAPF methods have been proposed to find collision-free paths given a representation of the $\mathcal{C}_{space}$. M* [3] conducts an A*-like search using a technique called subdimensional expansion where optimal individual paths for each robot are followed until a conflict occurs. Then, the conflicting agents are merged together and a new conflict-free composite path is found. Subdimensional Expansion RRT (sRRT) [19] applies subdimensional expansion to MRMP. It constructs individual policies by growing RRTs in the $C_{space}$ of each robot backward from the goal. It then expands a tree forward in the composite space by following the individual policies until conflict occurs. Then, new configurations are sampled for expansion in the composite space of the conflicting robots. While these methods perform well in open environments, in constrained settings, they expend excess computation on conflict resolution, limiting planning efficiency. Adaptive Robot Coordination (ARC) [20] first plans decoupled paths for each robot and then resolves conflicts by solving local subproblems in the composite spaces of the affected robots. Though it does focus computation within constrained settings, it does so reactively and spends excess time adjusting the local problem size to find the right resolution.
Kinodynamic Conflict-Based Search (K-CBS) [21] extends hybrid MRMP algorithms to accommodate kinodynamic constraints, drawing on the concepts of CBS-MP [2]. This approach introduces kinodynamic conflicts and employs a kinodynamic low-level planner akin to the kinodynamic RRT in [15], enabling the ability to solve high-dimensional problems with general dynamics.
@@ II-C Guided Motion Planning
Guided motion planning methods leverage external information to efficiently find paths. Prior work has explored using knowledge of the workspace in the form of a wavefront expansion [9] or workspace decomposition regions [10] to guide single-robot planning. Many approaches use workspace skeletons, embedded graphs in which edges represent free areas of the workspace and vertices represent connections between them. Examples include medial axis skeletons [22] for 2D environments and mean curvature skeletons [23] for 3D. They are generally quick to compute, for example, medial axis skeletons can be computed in $O(n\log n)$ time where $n$ is the number of obstacle edges [24]. Skeleton-guided motion planning methods [7, 8] use dynamic sampling regions, bounded local areas of the workspace that advance along skeleton edges, to efficiently explore free areas of the environment, including narrow passages, until a path is found.
@@
Prior work extends skeleton guidance to MRMP for scenarios with narrow passages which require a high level of coordination during planning. Composite Dynamic Region-biased RRT (CDR-RRT) [6] is a coupled method that lazily builds and searches over a composite skeleton, which is the Cartesian product graph of the skeleton for each robot in the group. MAPF is used to generate a path over the skeleton for each robot, taking into account each edge’s capacity as given by the clearance to obstacles in the environment. These paths are combined to form a path through the composite skeleton. Dynamic sampling regions advance over the composite skeleton edges, growing a tree until the goal is reached. Although CDR-RRT shows a significant improvement in scalability compared to other state-of-the-art methods in environments with narrow passages where coordination between the full robot group is required [6], considering the full composite space of all robots incurs significant computational overhead. We propose a hybrid method extended from CDR-RRT that uses topological guidance to inform when planning in the composite space (of all or a subgroup of robots) is necessary.
III Guided-DaSH
In this section, we describe our adaptation to the DaSH framework, Guided-DaSH, and explain the modifications we make to support motion planning problems without a highly structured search space and kinodynamic constraints. For a detailed description of the DaSH framework, see [5]. Figure 1 shows a comparison of Guided-DaSH and the original DaSH framework which we will elaborate on below. Algorithm 1 shows an overview of our framework.
Algorithm 1 Guided-DaSH
1:Heuristic $H$, Query $Q$
2:$\Pi\leftarrow\emptyset$triangleright Robot motion paths
3:$C\leftarrow\emptyset$triangleright Heuristic search constraints
@@ III Guided-DaSH
5:    $\mathcal{H}_{ts}\leftarrow\textsc{ConstructTaskSpaceHypergraph}(H,C,Q)$
6:    $\mathcal{X}_{ts}\leftarrow\textsc{OrderByDependency}(\mathcal{H}_{ts})$
7:    $\mathcal{H}_{m}\leftarrow\emptyset$triangleright Initialize motion hypergraph
8:    for $x_{ts}\in\mathcal{X}_{ts}$
9:         $\pi\leftarrow\textsc{ComputePath}(G,x_{ts})$
10:         if $\pi=\emptyset$
11:             $C\leftarrow C\cup\textsc{GetConstraint}(x_{ts})$
12:             break $\triangleright$ Restart heuristic search
13:         $\mathcal{X}_{conf}\leftarrow\textsc{FindCollisions}(\mathcal{H}_{m},\pi)$
14:         if $\mathcal{X}_{conf}\neq\emptyset$
15:             $C\leftarrow C\cup\textsc{GetConstraints}(\mathcal{X}_{conf}\cup\{x_{ts}\})$
16:             break $\triangleright$ Restart heuristic search
17:         $\mathcal{H}_{m}.\textsc{AddPath}(\pi)$
18:    $\Pi\leftarrow\textsc{ExtractMotionPaths}(\mathcal{H}_{m})$
III-A Task Space Hypergraph Construction
In the task space hypergraph $\mathcal{H}_{ts}=\{\mathcal{V}_{ts},\mathcal{E}_{ts}\}$, vertices $\mathcal{V}_{ts}$ are task space elements which represent local planning spaces and are composed of the robots required to coordinate at that point and relevant constraints they should adhere to. Hyperarcs $\mathcal{E}_{ts}$ can be categorized as composition, representing changes in the robots considered and constraints, or transition, representing changes in constraints for a fixed set of robots. The original DaSH framework [5] exhaustively constructs $\mathcal{H}_{ts}$; however, by using guidance to inform which task space elements are needed, we can efficiently construct a minimal portion of $\mathcal{H}_{ts}$. This makes our adapted framework advantageous for problems where exhaustive construction is intractable.
To construct a minimal portion of the task space hypergraph, we require guidance in the form of domain knowledge about the problem being considered. This defines the set of constraints which can exist in a task space element. Additionally, this knowledge must inform a heuristic $H$ required to search over the implicit problem space and construct a series of task space elements $\mathcal{V}_{ts}$ that satisfy a given query. Hyperarcs $\mathcal{E}_{ts}$ are added between subsequent task space elements to form $\mathcal{H}_{ts}$. The heuristic must also take as input a set of constraints that represent prohibited task space elements that cannot exist in the solution since they have failed to be translated into a motion solution. Searching over the implicit problem space can be achieved by searching over a smaller representative space that captures when coordination between robots is required. We provide an example in which information about the workspace topology is used to guide composition decisions in Section IV.
III-B Motion Hypergraph Construction
The motion hypergraph $\mathcal{H}_{m}$ stores the configurations and motions between them that make up the robots’ motion paths through the task space elements and hyperarcs from $\mathcal{H}_{ts}$. It is often necessary to leverage a guided planning method to ensure that the constraints associated with each task space element are met when constructing local paths. We detail an example of such a method in Section IV.
For each task space element and transition hyperarc in $x_{ts}\in\mathcal{H}_{ts}$, ordered from start to goal, a local path is computed in the composite space of the robots contained therein. These paths become vertices and motion hyperarcs in $\mathcal{H}_{m}$. Transitions often require a different method of path construction to account for, e.g., grasp constraints or other problem-specific factors. It is possible that constructing a local path for a task space element or transition could fail. In this case of a task space element failing, a constraint is added to the heuristic search to prohibit that task space element from existing in $\mathcal{H}_{ts}$. Similarly, in the case of a transition failing, constraints are added to prohibit the task space elements in the corresponding incoming and outgoing sets of the transition from existing simultaneously in $\mathcal{H}_{ts}$.
In the original DaSH formulation, the entire motion hypergraph is constructed before a scheduled query is performed that finds and resolves conflicts between robot paths through different task space elements or transitions. In our adapted framework; however, since we know from the heuristic search which task space element’s motions should occur concurrently, we can interleave motion hypergraph construction and conflict detection. By finding conflicts as they occur rather than after the motion hypergraph is constructed, we minimize wasted computation and can directly extract the motion solution from the motion hypergraph, bypassing the query stage.
Our modification to the conflict resolution stage also allows us to support kinodynamic planning. The scheduled query proposed in the original DaSH framework resolves conflicts by inducing waiting to prevent robots from colliding with each other. When planning under kinodynamic constraints, we cannot assume the ability of robots to instantaneously change velocity. Additionally, to avoid solving computationally difficult optimal control point-to-point problems, we cannot reuse local paths for task space elements occurring after conflicts in the robots’ paths. Thus, we must replan affected paths to avoid conflicts.
As we construct $\mathcal{H}_{m}$, after each local path is computed, we check for collisions with existing local paths occurring simultaneously by considering intermediate configurations along the paths. When a conflict is found, constraints are added to the heuristic search to prevent the corresponding task space elements or transitions from existing in $\mathcal{H}_{ts}$ together. When constraints are added, the heuristic search is redone, $\mathcal{H}_{ts}$ is recomputed, and construction of $\mathcal{H}_{m}$ restarts. This process continues until a conflict-free motion solution is found.
IV Workspace Guided-DaSH

### v4

# Scalable Multi-Robot Motion Planning Using Workspace Guidance-Informed Hypergraphs

@@ Abstract
In this work, we propose a method for multiple mobile robot motion planning that efficiently plans for robot teams up to 128 robots (an order of magnitude larger than existing state-of-the-art methods) in congested settings with narrow passages in the environment. We achieve this improvement in scalability by extending the state-of-the-art Decomposable State Space Hypergraph (DaSH) multi-robot planning framework to support mobile robot motion planning in congested environments. This is a problem that DaSH cannot be directly applied to because it lacks a highly structured, easily discretizable task space and features kinodynamic constraints. We accomplish this by exploiting knowledge about the workspace topology to limit exploration of the planning space and through modifying DaSH’s conflict resolution scheme. This guidance captures when coordination between robots is necessary, allowing us to decompose the intractably large multi-robot search space while limiting risk of inter-robot conflicts by composing relevant robot groups together while planning.
@@ I Introduction
Multi-robot systems are becoming increasingly prevalent in settings like warehouses and factories. These often constrained environments require careful yet computationally fast motion planning for each robot to accomplish its task without collision. Some navigation approaches compute individual decoupled paths for each robot and resolve conflicts online [sl-uppccdpmrs-2002]; however, in congested environments, this introduces the risk of deadlock. Alternatively, coupled offline multi-robot motion planning (MRMP) approaches [sl-uppccdpmrs-2002] plan for all robots together in the composite space, the joint planning space of all robots. While these methods provide the high level of coordination required to solve difficult problems, due to the large size of the search space, they are computationally intractable for large robot teams. Hybrid methods [smsa-romrmpucbs-21, wc-mcmppapb-2011, hpksga-tpqs-2018] seek to combine the benefits of decoupled and coupled planning, providing an increased level of coordination when necessary while also maintaining scalability. While existing approaches [smqma-arc-24, wkc-pppfmrwse-12] reactively provide coordination after conflicts occur during planning, our proposed method aims to improve efficiency by predicting when coordination between robots is necessary.
To model when coordination between robots is needed, we build our method off of the Decomposable State Space Hypergraph (DaSH) [mcbma-hbmrtamp-23] framework. DaSH presents a general structure for hybrid planning that benefits from a sparse high-level representation of the task space that composes robots into the same planning space when necessary and allows the planner to stay in low-dimensional search spaces when possible. This representation directs the construction of a motion solution. DaSH proposes an approach to exhaustively constructing the high-level representation that performs well for problems like multi-manipulator rearrangement [mcbma-hbmrtamp-23], where the task space has an exploitable structure since actions can be classified as pick, place, handoff, etc., but becomes intractable for problems which lack this highly structured task space. Because of this, we cannot directly apply DaSH to mobile robot motion planning.
Figure 1: A comparison of the DaSH framework and Workspace Guided-DaSH showing an example task space hypergraph for each. Hypergraph vertices show which robots and/objects are present in that task space element. (left) DaSH’s exhaustive construction of the task space hypergraph for a multi-manipulator rearrangement with two robots and one object. (right) Our method uses workspace guidance to induce structure onto problems that lack this natural structure, which makes construction of a minimal portion of the task space hypergraph possible. Here, robots are composed into the same planning space based on their movement along a topological skeleton (details in Sec. III).
@@
Here, we propose extending DaSH for mobile robot motion planning in environments with narrow passages, a classical challenge in motion planning exacerbated in the multi-robot setting, where the robots may also have kinodynamic constraints that further complicate planning. Our insight is that we can leverage information about the planning problem to induce this structure while building the high-level representation and further exploit this guidance, in conjunction with a modified conflict resolution scheme, to efficiently search the planning space and support kinodynamic constraints. In mobile robot motion planning, we can effectively use workspace information to provide this guidance [mmuma-smrmpcetg-23, dsba-drbrrt-16, suda-tgrcdrs-20, rsb-beaeisbmp-2014, pkv-mpdsclp-10]. As such, we extend the DaSH framework to present Workspace Guided-DaSH, a novel method for the problem of multiple mobile robot motion planning in environments with narrow passages. Our contributions include:
@@
  • A method that extends the DaSH framework to a multi-robot planning problem without a highly structured task space by leveraging the workspace to guide planning in multiple ways.
@@
  • Support for planning under kinodynamic constraints via an adapted conflict resolution structure.
@@ II-A Motion Planning Preliminaries
In motion planning, a configuration refers to a set of values for a robot’s degrees of freedom, including their position in the workspace, orientation, and other parameters such as joint angles. The configuration space ($\mathcal{C}_{space}$) consists of all possible configurations and can be partitioned into $\mathcal{C}_{free}$ and $\mathcal{C}_{obst}$ made up of valid and invalid configurations, respectively. Motion planning is the problem of finding a path from a start configuration $q_{s}$ to a goal configuration $q_{g}$ through $\mathcal{C}_{free}$. Computing the full $\mathcal{C}_{obst}$ is intractable [ss-otpmpiigtfctporam-83, c-crmp-88], thus state-of-the-art methods turned to sampling-based motion planning [kslo-prpp-96, l-rrtntpp-1998]. Two commonly used methods, Rapidly-exploring Random Trees (RRT) [l-rrtntpp-1998] and Probabilistic Roadmaps (PRM) [kslo-prpp-96] build tree-based and graph-based representations of $\mathcal{C}_{free}$ respectively. These methods have also been extended to problems that require satisfying kinematic and dynamic constraints. Kinodynamic RRT [hklr-rkmpwmo-02] extends RRT to work with kinodynamic constraints through extensions of randomly sampled control inputs for random durations until a goal region is reached.
@@ II-B Multi-robot Motion Planning
The MRMP problem is an extension of the motion planning problem that consists of finding valid paths $\Pi=\{\pi_{r_{1}},...,\pi_{r_{n}}\}$ for a set of robots $R=\{r_{1},...,r_{n}\}$ that also avoid inter-robot collision. Multi-agent pathfinding (MAPF) is the discrete state space equivalent of the MRMP problem, which consists of finding collision-free paths for a set of robots over a given graph representation (e.g., a grid) rather than through continuous space. In environments with narrow passages, it is non-trivial to construct a graph representation that captures the connectivity of the planning space, hence our consideration of MRMP rather than MAPF.
Some coupled methods, including Composite PRM [sl-uppccdpmrs-2002] and Composite RRT [l-rrtntpp-1998], provide a high level of coordination by planning directly in the composite space, $C_{0}\times...\times C_{n-1}$ where $C_{i}$ is the configuration space of robot $i$. Other coupled methods, including MRdRRT [ssh-faniaehdrfeoirimm-16] and its variants [dsshb-sammp-17, ssdhb-dsaiammp-20], construct single-robot roadmaps and then use these to search an implicit composite space roadmap. Searching the composite space, even implicitly, becomes intractable with large groups of robots due to the high dimensionality of the space.
Hybrid methods alternate between planning in spaces with different compositions of robots. CBS-MP [smsa-romrmpucbs-21] performs a low-level search to find decoupled paths for each robot and resolves conflicts using a high-level search in the composite space. Kinodynamic Conflict-Based Search (K-CBS) [kal-cbsfmrmpwkc-22] extends hybrid MRMP algorithms to accommodate kinodynamic constraints, drawing on the concepts of CBS-MP [smsa-romrmpucbs-21]. Subdimensional Expansion RRT (sRRT) [wkc-pppfmrwse-12] constructs individual policies by growing RRTs in the $\mathcal{C}_{space}$ of each robot backward from the goal. It then expands a tree forward in the composite space by following the individual policies until conflict occurs. Then, new configurations are sampled for expansion in the composite space of the conflicting robots. While these methods perform well in open environments, in constrained settings, they expend excess computation on conflict resolution, limiting planning efficiency. Adaptive Robot Coordination (ARC) [smqma-arc-24] plans decoupled paths and then resolves conflicts by solving local subproblems in the composite spaces of the affected robots. Though it does focus computation within constrained settings, it does so reactively and spends excess time adjusting the local problem size to find the right resolution. Its kinodynamic variant [qsmma-karcfmrkp-2025] similarly resolves collisions and ensures kinodynamic feasibility using trajectory optimization.
@@ II-C Guided Motion Planning
Guided motion planning methods use external information to efficiently find paths. Many single-robot approaches use workspace skeletons, embedded graphs in which edges represent free areas of the workspace and vertices are connections between them. Examples include medial axis skeletons [Blum_1967_6755] for 2D environments and mean curvature skeletons [t-mcs-12] for 3D. They are generally quick to compute, for example, medial axis skeletons can be computed in $O(n\log n)$ time where $n$ is the number of obstacle edges [l-matoaps-1982]. Skeleton-guided planning methods [dsba-drbrrt-16, suda-tgrcdrs-20] use dynamic sampling regions, bounded local areas of the workspace that advance along skeleton edges, to explore free areas of the environment, including narrow passages, until a path is found.
@@
Prior work extends skeleton guidance to MRMP for scenarios with narrow passages, which require a high level of coordination during planning. Composite Dynamic Region-biased RRT (CDR-RRT) [mmuma-smrmpcetg-23] is a coupled method that lazily builds and searches over a composite skeleton, which is the Cartesian product graph of the skeleton for each robot in the group. MAPF is used to generate a path over the skeleton for each robot, taking into account each edge’s capacity as given by the clearance to obstacles in the environment. These paths are combined to form a path through the composite skeleton. Dynamic sampling regions advance over the composite skeleton edges, growing a tree until the goal is reached. Although CDR-RRT shows a significant improvement in scalability compared to other state-of-the-art methods in environments with narrow passages where coordination between the full robot group is required [mmuma-smrmpcetg-23], considering the full composite space of all robots incurs significant computational overhead. We propose a hybrid method extended from CDR-RRT that uses topological guidance to inform when planning in the composite space (of all or a subgroup of robots) is necessary.
II-D Decomposable State Space Hypergraph Framework
The Decomposable State Space Hypergraph (DaSH) framework [mcbma-hbmrtamp-23] for hybrid multi-robot planning leverages a directed hypergraph representation to enable coordination between relevant groups of robots without considering the full composite space unless necessary. A directed hypergraph $\mathcal{H}=(\mathcal{V},\mathcal{E})$ is a generalization of a graph where hyperarcs $E\in\mathcal{E}$ represent directed connections between sets of vertices. This representation provides the advantage of sparsity. While an equivalent graph could be constructed, vertices would need to represent a state for every robot in the problem, leading to a combinatorial explosion as the number of robots increases. The hypergraph representation, however, only includes relevant robots in each vertex or hyperarc, allowing groups of robots to be composed and decomposed. This is especially beneficial for sampling-based motion planning where the planning time is impacted by the dimension of the $\mathcal{C}_{space}$.
The DaSH framework consists of three components: construction of the task space hypergraph, construction of the motion hypergraph, and the query. The task space hypergraph, $\mathcal{H}_{ts}$, is a high-level representation of the task space that captures when coordination between robots is required. allowing robots to move between planning spaces with different compositions of robots at different stages of planning. In DaSH, these planning spaces, which are represented by the vertices of the task space hypergraph, are referred to as task space elements. Task space elements are made up of a set of relevant robots and constraints (e.g., start/goal constraints, path constraints). The DaSH framework has been implemented for problems including multi-manipulator rearrangement [mcbma-hbmrtamp-23] where the task space is highly structured and easily discretizable with task space elements representing modes such as a robot holding an object. This structure allows the sparse task space hypergraph to be constructed exhaustively. For task spaces without such structure, however, this exhaustive construction becomes intractable.
The task space hypergraph is used to build a low-level motion hypergraph, $\mathcal{H}_{m}$, encoding motion paths. Here, vertices represent configurations and hyperarcs represent local paths between them. The query phase consists of two parts. First, an optimistic solution on the motion hypergraph is found which ensures a valid transition history (i.e., guaranteeing that each robot exists in exactly one place at each point along the solution). Second, because hyperarcs in the motion hypergraph give a local plan for the subset of robots in the group being considered and do not make assumptions about the locations of the other robots, a scheduled query must then be used to validate the solution in terms of inter-robot collisions by inducing waiting.
Considering extending DaSH to a problem without a naturally highly structured task space, each of these components poses a design question, which we discuss in Section III. Additionally, we discuss the modifications to the conflict resolution structure that we make to accommodate kinodynamic planning since we can no longer assume the robots are able to instantaneously change velocity in the scheduled query. These changes allow us to bypass the query stage entirely.
Algorithm 1 Workspace Guided-DaSH
1:Workspace Skeleton $S$, Query $Q$
[…]
@@ II-D Decomposable State Space Hypergraph Framework
5:  $\Pi^{S}\leftarrow\textsc{MAPF}(S,C,Q)$ $\triangleright$ Find paths over the skeleton
6:  $\mathcal{H}_{ts}\leftarrow\textsc{ConstructTaskSpaceHypergraph}(\Pi^{S})$
7:  $\mathcal{X}_{ts}\leftarrow\textsc{OrderByDependency}(\mathcal{H}_{ts})$
8:  $\mathcal{H}_{m}\leftarrow\emptyset$ $\triangleright$ Initialize motion hypergraph
9:  for $x_{ts}\in\mathcal{X}_{ts}$ $\triangleright$ Iterate over $\mathcal{V}_{ts}$ and $\mathcal{E}_{ts}$
10:    $\pi\leftarrow\textsc{ComputePath}(G,x_{ts})$
11:    if $\pi=\emptyset$ $\triangleright$ Failed to find a path
12:     $C\leftarrow C\cup\textsc{GetConstraint}(x_{ts})$
13:     break $\triangleright$ Restart heuristic search
14:    $\mathcal{X}_{conf}\leftarrow\textsc{FindCollisions}(\mathcal{H}_{m},\pi)$
15:    if $\mathcal{X}_{conf}\neq\emptyset$ $\triangleright$ Found conflict
16:     $C\leftarrow C\cup\textsc{GetConstraints}(\mathcal{X}_{conf}\cup\{x_{ts}\})$
17:     break $\triangleright$ Restart heuristic search
18:    $\mathcal{H}_{m}.\textsc{AddPath}(\pi)$
19:  $\Pi\leftarrow\textsc{ExtractMotionPaths}(\mathcal{H}_{m})$
III Workspace Guided-DaSH Method